Solve the equation by factoring.
2 - 10x = (3x - 7)(x + 1)

Answer by ankor@dixie-net.com(22740)

Solve the equation by factoring.
2 - 10x = (3x - 7)(x + 1)
FOIL
2 - 10x = 3x^2 + 3x - 7x - 7
combine like terms on the right
0 = 3x^2 + 3x - 7x + 10x - 7 - 2
0 = 3x^2 + 6x - 9
simplify, divide by 3
x^2 + 2x - 3 = 0
Factors to
(x+3)(x-1) = 0
x = -3
and
x = 1